OpenStudy (studyx3):

Medal and fan!!!!! A culture started with 2,000 bacteria. After 8 hours, it grew to 2,400 bacteria. Predict how much bacteria will be present after 19 hours. Round answer to the nearest whole number.

OpenStudy (thekingofcr):

Because the bacteria culture grew by 200 bacteria in 6 hours, the unit rate of that is 33 1/3 bacteria growing per hour. If you multiply that by 12, you get 400. If you add 2,000 + 400 = 2,400 bacteria growing

OpenStudy (studyx3):

Thank you!

OpenStudy (thekingofcr):

2633.33333333 You can search for any mathematical expression, using functions such as: sin, cos, sqrt, etc. You can find a complete list of functions here. RadDeg x! Inv sin ln π cos log e tan √ Ans EXP xy ( ) % AC 7 8 9 ÷ 4 5 6 × 1 2 3 − 0 . = +

OpenStudy (thekingofcr):

final answer 2633.33333333

OpenStudy (thekingofcr):

im pretty sure lol

OpenStudy (wolf1728):

First we solve for the rate k = (ln [y(t) ÷ a]) ÷ t where y(t) is ending amount a is beginning amount t is time k = (ln (2400/2000)) / 8 k = (ln (6/5)) / 8 k = ln(1.2) / 8 k = 0.18232155679 / 8 k = 0.0227901946 y(t) = a × e k*t endg amt = 2,000 x 2.718281828^(0.0227901946 * 19) endg amt = 2,000 x 2.718281828^0.4330136974 endg amt = 2,000 x 1.5418973404 endg amt = 3,084